Consider a monopolist facing the demand curve p = 90 − 2q with cost function c(q) = 0.25q^2 .

(a) Find the profit-maximizing quantity qm and price pm. What are the monopolist’s profits?

(b) What is the value of the Lerner index at qm?

(c) Find the efficient quantity and draw a graph depicting the deadweight loss under monopoly.
Answer #1

P-90 - 29 TR= P2 = 909 - 29 MR= 90 - 49 ( = 0.2592 mc= 0.59 MR=mc 90-49=0.58 920 Pm = 90-2 (20) - So Profit a so(20) - 0.25(2